Two men are diametrically on the opposite sides of a tower.They measure the angle of diameter of the top of a tower as 27 degree and 34 degree respectively.If the height of the tower is 50m.Find the distance between the two men.
Answer by nyc_function(2741)

Let d1 = distance 1
Let d2 = distance 2
let TD = total distance between the two men
tan(27) = 50/d1 and tan(34) = 50/d2
So d1 = 50/tan(27) and d2 = 50/tan(34)
d1 + d2 = 98.13 + 74.13
TD = 172.26 meters